my problem is : The length of a rectangular-shaped backyard is 8 feet less than twice the width. If 260 feet of fencing is needed to enclose the yard, find the dimensions of the yard.
Answer by nerdybill(7384)

The length of a rectangular-shaped backyard is 8 feet less than twice the width. If 260 feet of fencing is needed to enclose the yard, find the dimensions of the yard.
.
Let w = width
then from "length of a rectangular-shaped backyard is 8 feet less than twice the width" we have:
2w-8 = length
.
perimeter = 2(width + length)
260 = 2(w + 2w-8)
260 = 2(3w-8)
130 = 3w-8
138 = 3w
46 feet = w (width)
.
length:
2w-8 = 2(46)-8 = 92-8 = 84 feet (length)